OpenStudy (anonymous):

Solve for theta: Giving a general formula for all of the solution tan(theta)=-1

OpenStudy (anonymous):

\[\tan(\theta)=-1\] \[\theta =\frac{3\pi}{4} \] is one solution

OpenStudy (apoorvk):

tan theta = -1 theta = n*pi + 3pi/4, where n is any integer (the general solution)

OpenStudy (anonymous):

since tangent is periodic with period pi, all the solutions will look like \[\frac{\pi}{4}+k\pi, k \in \mathbb Z\]

OpenStudy (anonymous):

typo there i meant \[\frac{3\pi}{4}+k\pi, k \in \mathbb Z\]